OpenStudy (anonymous):

Solve by substitution method: 7m +n =46 m-3n=16

OpenStudy (anonymous):

7m + n = 46 ----(x) m - 3n = 16 ----(xx) multiply eq (xx) by 7 and subtract it from (x). 7m + n = 46 7m - 21n = 112 -_____________ 22n = -66 n = -66/22 = -3 n = -3 put the value of n in any of the above eq say (x) 7m + (-3) = 46 7m - 3 = 46 7m = 46 + 3 = 49 m = 49/7 =7 m = 7